OpenStudy (anonymous):

3x-4y=12 i need help finding the slope and y intercept

OpenStudy (amistre64):

what methods have you tried so far?

OpenStudy (anonymous):

-4y = -3x + 12 y = (3/4)x - 3 Slope = (3/4) YIntercept = (0, -3)

OpenStudy (anonymous):

thank you!